# DAISI: Data Assimilation with Inverse Sampling using Stochastic Interpolants

### v1

@@ Abstract
Data assimilation (DA) is a cornerstone of scientific and engineering applications, combining model forecasts with sparse and noisy observations to estimate latent system states. Classical DA methods, such as the ensemble Kalman filter, rely on Gaussian approximations and heuristic tuning (e.g., inflation and localization) to scale to high dimensions. While often successful, these approximations can make the methods unstable or inaccurate when the underlying distributions of states and observations depart significantly from Gaussianity. To address this limitation, we introduce DAISI, a scalable filtering algorithm built on flow-based generative models that enables flexible probabilistic inference using data-driven priors. The core idea is to use a stationary, pre-trained generative prior to assimilate observations via guidance-based conditional sampling while incorporating forecast information through a novel inverse-sampling step. This step maps the forecast ensemble into a latent space to provide initial conditions for the conditional sampling, allowing us to encode model dynamics into the DA pipeline without having to retrain or fine-tune the generative prior at each assimilation step. Experiments on challenging nonlinear systems show that DAISI achieves accurate filtering results in regimes with sparse, noisy, and nonlinear observations where traditional methods struggle.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Our method, DAISI, performs data assimilation by combining a flow-based unconditional prior $\mathbb{P}_{\infty}$ with the forecast ensemble $\hat{\pi}_{n}$. To condition on an observation ${\bm{y}}_{n}$, one could apply the guided SDE starting from random noise, producing $\mathbb{P}_{\infty}^{{\bm{y}}}$ and ignoring the forecast (blue). Instead, DAISI (green): (i) applies the backward SDE to the forecast ensemble, producing inverted samples $\Psi_{\sharp}\hat{\pi}_{n}$, and (ii) uses these latents as initial conditions for the guided SDE, generating approximate samples from the filtering distribution $\pi_{n}$.
@@
Estimating the evolving state of a complex dynamical system is a fundamental challenge in science and engineering. In many real-world problems—such as weather forecasting, fluid mechanics, neuroscience and robotics—the states of interest are only partially observed and subject to multiple sources of uncertainty (asch_DA_2016). DA, and in particular, filtering, seeks to combine imperfect model forecasts with sparse, noisy observations to reconstruct these hidden states. Mathematically, this corresponds to estimating the filtering distribution $p({\bm{x}}_{n}|{\bm{y}}_{1:n})$ where ${\bm{x}}_{n}$ denotes the latent state and ${\bm{y}}_{1:n}:=({\bm{y}}_{1},\,\dots,\,{\bm{y}}_{n})$, where ${\bm{y}}_{n}$ is the observation at time $n$ generated via the likelihood $p({\bm{y}}_{n}|{\bm{x}}_{n})$.
@@
Weather forecasting provides a clear illustration of the challenges present in DA. The atmosphere is chaotic, high-dimensional, and observed through nonlinear, noisy measurements. Accurate state estimation is essential for applications ranging from early warning systems and agriculture to aviation safety and renewable energy management. (noaa2025billiondollar; whitt2023economic; ipcc2023synthesis; energy_pred). To meet these demands, classical DA methods, such as the Ensemble Kalman Filter (EnKF), or variational methods such as 4DVar, have been developed over decades of DA research. However, these have respective drawbacks; EnKF is only guaranteed to work under near-Gaussian settings (calvello2024accuracy), and the inflation and localization parameters necessary to stabilize the filter are notoriously challenging to tune (bannister2017review). 4DVar, on the other hand, requires the tedious development of an adjoint model and cannot quantify uncertainty, due to it being a MAP estimation method. Particle filters (naesseth2019elements) can, in principle, solve the filtering problem, but suffers from the curse of dimensionality (bengtsson2008curse). These limitations motivate the development of more flexible, data-driven approaches to DA.
@@
Recent progress in offline inverse problems—such as those in medical imaging, astrophysics, and computer vision—has shown that flow- and diffusion-based generative models can serve as powerful priors for conditional sampling (zhenginversebench). However, extending these ideas to sequential filtering presents unique challenges. One possibility is to learn a generative prior for the predictive distribution $p({\bm{x}}_{n}|{\bm{y}}_{1:n-1})$, but this would require retraining the model at every time step $n$ (bao_1_trained) , making it impractical for operational use. Another strategy is to use a pre-trained generative forecast model to guide towards observations (chen2025flowdas; savary2025training), though such models require the use of specialized dynamical models. Alternatively, approaches based on approximating the smoothing distribution $p({\bm{x}}_{0:n}|{\bm{y}}_{1:n})$ (rozet2023score) can, in principle, be used for filtering. However, they are memory-intensive and scale poorly with the number of time steps. We elaborate on how these approaches relate to our proposed method in Section 3.3.
@@ 1.1 Contributions
In this work, we present DAISI (Data Assimilation with Inverse sampling using Stochastic Interpolants), a scalable filtering algorithm built on flow-based generative models. DAISI avoids retraining a generative model at each assimilation step by leveraging a stationary pre-trained generative prior to condition on new observations via guidance. To integrate dynamical information, DAISI couples this generative prior with a forecast model that advances an ensemble of states in time. Following the forecast, an inverse-sampling step runs the generative SDE backward from the forecast ensemble, mapping forecasted states to latent variables. These latent states then serve as initial conditions for conditional sampling under the learned prior. The full procedure is illustrated in Figure 1.
@@
    Modular design, supporting any flow-based generative model or gradient-based guidance method.
@@
    Expressive uncertainty quantification, capturing complex, multimodal posteriors under sparse, noisy, and nonlinear observations.
@@ 2.1 Data Assimilation
where $\mathcal{F}_{\sharp}\pi$ denotes the pushforward of the measure $\pi$ with respect to a map $\mathcal{F}$, defined by $\mathcal{F}_{\sharp}\pi:=\mathrm{Law}(\mathcal{F}(X))$ for $X\in\pi$. Starting from $\pi_{0}$, repeated applications of (3)–(4) yields the sequence $\pi_{1},\ldots,\pi_{N}$. Although this recursion is well defined, computing these measures exactly is typically infeasible in practice. Our aim is therefore to develop a practical and scalable approximation to the cycle (3)–(4), capable of operating reliably in high-dimensional settings with sparse, nonlinear observations.
@@ 2.2 Stochastic Interpolants for Generative Modeling
Our proposed DA method leverages recent advances in generative modeling. In the static setting (i.e., ignoring time evolution), flow-based models enable sampling from a target distribution $\rho_{1}$ by learning a transport map from a simple latent distribution $\rho_{0}$. Such transports can be parameterized in multiple ways; in this work, we adopt the framework of stochastic interpolants (albergo2023stochastic), though closely related formulations appear in the probability flow ODE for diffusion models (song2020score) and in flow matching (lipmanflow).
@@
Let $\rho_{t}:=\mathrm{Law}({\bm{z}}_{t})$ denote the distribution of ${\bm{z}}_{t}$ with density $p_{t}$, and define its score ${\bm{s}}(t,{\bm{z}}):=\nabla\log p_{t}({\bm{z}})$ and drift ${\bm{b}}(t,{\bm{z}}):=\mathbb{E}[\dot{{\bm{z}}}_{t}|{\bm{z}}_{t}={\bm{z}}]$. In albergo2023stochastic, it is shown that for any non-negative $\epsilon\in C^{0}([0,1])$, the SDE
@@ 2.3 Conditional Generation via Guidance
In general, the guidance term is intractable as we usually only have access to $p({\bm{y}}|{\bm{z}}_{1})$ but not $p({\bm{y}}|{\bm{z}}_{t})=\mathbb{E}_{{\bm{z}}_{1}|{\bm{z}}_{t}}[p({\bm{y}}|{\bm{z}}_{1})]$. Various approaches exist to approximate this term, e.g. diffusion posterior sampling (DPS) (chung2023diffusion) and moment matching posterior sampling (MMPS) (Rozet_Andry_Lanusse_Louppe_2024) (see Appendix A.5 for details and (daras2024survey) for a survey on guidance methods).
@@ 3 Method
Our method, DAISI, performs filtering by combining the strengths of ensemble-based methods with the expressive priors offered by flow-based generative models. This is conceptually similar to the hybrid ensemble-variational (EnVar) method in classical DA (hamill2000hybrid; lorenc2003potential), which blend a static climatological background covariance $\mathbf{P}$ with an “error-of-the-day” ensemble covariance to represent forecast uncertainty. This compensates for the limitations of either component alone: relying solely on a static background covariance ignores the dynamical evolution of errors, while using only a small ensemble fails to characterize high-dimensional uncertainties accurately.
DAISI builds on this philosophy but moves beyond Gaussian background covariances $\mathbf{P}$ used in hybrid EnVar by considering a full background measure $\mathbb{P}_{\infty}$, learned using flow-based generative models. In practice, we take $\mathbb{P}_{\infty}$ to be the invariant measure of the dynamical system (1), whose samples can be generated via long-time rollouts. While its existence and ergodicity are not generally guaranteed, it is a reasonable assumption for many geophysical flows, which are typically dissipative and subject to stochastic forcing. Given the learned prior $\mathbb{P}_{\infty}$, we can then assimilate new observations ${\bm{y}}_{n}$ by sampling from the posterior
@@
To achieve this, we introduce inverse sampling, which transfers dynamical information from the ensemble forecast into the latent space of the generative model. Given forecast particles $\{\hat{{\bm{x}}}^{(j)}_{n}\}_{j=1}^{J}$, this proceeds by solving the unconditional backward SDE (7) from $t=1$ to $t=t_{\min}\in(0,1)$, using each $\hat{{\bm{x}}}_{n}^{(j)}$ as terminal conditions. The resulting latent variables $\{{\bm{z}}_{t_{\min},n}^{(j)}\}_{j=1}^{J}$ encode the forecast information in “noise space”. We can then perform conditional sampling by integrating the guided forward SDE (8) from $t_{\min}$ to 1, initialized at these latent states. This produces updated particles $\{{\bm{x}}_{n}^{(j)}\}_{j=1}^{N}$ that are samples of $\mathbb{P}_{\infty}^{\bm{y}}$, while simultaneously containing information about the ensemble $\{\hat{{\bm{x}}}^{(j)}_{n}\}_{j=1}^{J}$ through its latent representations $\{{\bm{z}}_{t_{\min},n}^{(j)}\}_{j=1}^{J}$.
@@ 3.1 Does DAISI sample from the filtering distribution?
where $\hat{\pi}_{n}$ denotes the predictive distribution at time $n$, and the expectation is over the Brownian motions used when integrating the backward and forward SDEs. For DAISI to function as a reliable filter, this measure should approximate the true filtering distribution $\pi_{n}$ closely. While there is no reason to believe that (13) matches $\pi_{n}$ exactly, we argue that we can achieve a reasonably good match by tuning the hyperparameters $t_{\min}$ and $\epsilon$ appropriately.
@@
We first examine the limiting cases $t_{\min}\rightarrow 1$ and $t_{\min}\rightarrow 0$, assuming $\epsilon=0$ (deterministic flows). By time-continuity of the flows on $[0,1]$, the first limit $t_{\min}\rightarrow 1$ trivially yields $\pi_{n,t_{\min},0}^{\text{DAISI}}\rightarrow\hat{\pi}_{n}\propto f_{n}\mathbb{P}_{\infty}$. Comparing with (16), we observe that although the density ratios agree, the underlying base measures differ.
@@
$\widetilde{\rho}_{0}$ $:=(\Psi_{1,0}^{0})_{\sharp}\hat{\pi}_{n}\propto(f_{n}\circ(\Psi_{1,0}^{0})^{-1})(\Psi_{1,0}^{0})_{\sharp}\mathbb{P}_{\infty}$
$=(f_{n}\circ(\Psi_{1,0}^{0})^{-1})\rho_{0},$ (17)
where we used that $(\Psi_{1,0}^{0})_{\sharp}\mathbb{P}_{\infty}=\rho_{0}$. Next applying the guided flow in the second step gives us
$\pi_{n,0,0}^{\text{DAISI}}=(\Phi^{{\bm{y}},0}_{0,1})_{\sharp}\widetilde{\rho}_{0}$ (18)
@@
where $g_{n}^{{\bm{y}}}:=f_{n}\circ(\Phi^{{\bm{y}},0}_{0,1}\circ\Psi^{0}_{1,0})^{-1}$ and we used that $\Phi^{{\bm{y}},0}_{0,1}$ transports $\rho_{0}$ to $\mathbb{P}_{\infty}^{\bm{y}}$ by construction. Comparing (19) with (16), we see that in this limit, the base measures match, while the density ratios differ. Thus, the extremes $t_{\min}\rightarrow 1$ and $t_{\min}\rightarrow 0$ each match only one component of the true filtering distribution, suggesting the existence of an intermediate $t_{\min}\in(0,1)$ that provides the best trade-off between these two types of mismatch.
In Figure 2, we illustrate this trade-off using a simple 1D toy problem (see Appendix D.2 for details). When $t_{\min}=0.01$, the distirbution $\pi^{\text{DAISI}}_{n}$ (in orange) shares characteristics of $\mathbb{P}^{\bm{y}}_{\infty}$, as predicted from (19). Conversely, for a large value such as $t_{\min}=0.6$, the distribution closely matches $\hat{\pi}_{n}$, as expected. The intermediate setting $t_{\min}=0.3$ yields a distribution that interpolates between these two extremes, yielding a distribution that better aligns with the filtering distribution, which combines features of both $\mathbb{P}^{\bm{y}}_{\infty}$ and $\hat{\pi}_{n}$ — as reflected by the noticeably lower Maximum Mean Discrepancy (MMD) score (gretton2012kernel).
@@
Assume that the marginal laws $\{\rho_{t}\}_{t\in[0,1]}$ and $\{\rho_{t}^{\bm{y}}\}_{t\in[0,1]}$ of the interpolants bridging $\rho_{0}=\mathcal{N}(\bm{0},\mathbf{I})$ with $\rho_{1}=\mathbb{P}_{\infty}$ and $\rho_{1}=\mathbb{P}_{\infty}^{\bm{y}}$, respectively, satisfy a log-Sobolev inequality with constant $\lambda>0$, uniform in $t$. Then,
@@
We see from this result that in the case $\epsilon=0$, the Kullback–Leibler divergence between the DAISI output and the guided reference measure is preserved, so the conditioning step retains all information carried by the forecast ensemble. On the other hand, when $\epsilon\rightarrow\infty$, we have $\pi_{n,0,\epsilon}^{\text{DAISI}}\rightarrow\mathbb{P}_{\infty}^{\bm{y}}$ and therefore all ensemble information is lost. In practice, we find that taking a small but non-zero $\epsilon$ is useful for mitigating the tendency of particles from collapsing to a single state as assimilation progresses. In our 1D toy experiment (Appendix D.2), we show that tuning both $t_{\min}$ and $\epsilon$ yields better performance than adjusting $t_{\min}$ alone.

### v2

@@ Abstract
Data assimilation (DA) is a cornerstone of scientific and engineering applications, combining model forecasts with sparse and noisy observations to estimate latent system states. Classical high-dimensional DA methods, such as the ensemble Kalman filter, rely on Gaussian approximations that are violated for complex dynamics or observation operators. To address this limitation, we introduce DAISI, a scalable filtering algorithm built on flow-based generative models that enables flexible probabilistic inference using data-driven priors. The core idea is to use a stationary, pre-trained generative prior that first incorporates forecast information through a novel inverse-sampling step, before assimilating observations via guidance-based conditional sampling. This allows us to leverage any forecasting model as part of the DA pipeline without having to retrain or fine-tune the generative prior at each assimilation step. Experiments on challenging nonlinear systems show that DAISI achieves accurate filtering results in regimes with sparse, noisy, and nonlinear observations where traditional methods struggle.
Machine Learning, ICML, Filtering, Diffusion models, Data Assimilation, Generative models
@@ 1 Introduction
Figure 1: DAISI combines a flow-based unconditional prior $\mathbb{P}_{\infty}$ with the forecast ensemble $\hat{\pi}_{n}$. To condition on an observation ${\bm{y}}_{n}$, one could apply the guided SDE starting from random noise, producing $\mathbb{P}_{\infty}^{{\bm{y}}}$ (blue). However, this would ignore the information contained in the forecast. Instead, DAISI (green): (i) applies the backward SDE to the forecast ensemble, producing inverted samples $\Psi_{\sharp}\hat{\pi}_{n}$, and (ii) uses these latents as initial conditions for the guided SDE, generating approximate samples from the filtering distribution $\pi_{n}$.
@@
Estimating the evolving state of a complex dynamical system is a fundamental challenge in science and engineering. In many real-world problems—such as weather forecasting, fluid mechanics, neuroscience and robotics—the states of interest are only partially observed and subject to multiple sources of uncertainty (Asch et al., 2016). DA, and in particular, filtering, seeks to combine imperfect model forecasts with sparse, noisy observations to reconstruct these hidden states. Mathematically, this corresponds to estimating the filtering distribution $p({\bm{x}}_{n}|{\bm{y}}_{1:n})$ where ${\bm{x}}_{n}$ denotes the latent state and ${\bm{y}}_{1:n}:=({\bm{y}}_{1},\,\dots,\,{\bm{y}}_{n})$, where ${\bm{y}}_{n}$ is the observation at time $n$ modeled via the likelihood $p({\bm{y}}_{n}|{\bm{x}}_{n})$.
@@
Weather forecasting provides a clear illustration of the challenges present in DA. The atmosphere is chaotic, high-dimensional, and observed through nonlinear, noisy measurements. Accurate state estimation is essential for applications ranging from early warning systems and agriculture to aviation safety and renewable energy management (NOAA NCEI, 2025; Whitt and Gordon, 2023; IPCC, 2023; Andrade and Bessa, 2017). To meet these demands, classical DA methods, such as the Ensemble Kalman Filter (EnKF), or variational methods such as 4DVar, have been developed over decades of DA research. However, these methods have respective drawbacks; EnKF is only guaranteed to work under near-Gaussian settings (Calvello et al., 2024), and the inflation and localization parameters necessary to stabilize the filter are notoriously challenging to tune (Bannister, 2017). 4DVar, on the other hand, requires the tedious development of an adjoint model and cannot quantify uncertainty, due to it being a MAP estimation method. These methods have also been shown to struggle when applied to ML-based forecasting models (Tian et al., 2024). Particle filters (e.g., Naesseth et al. (2019)) can, in principle, solve the filtering problem, but suffers from the curse of dimensionality (Bengtsson et al., 2008). These limitations motivate the development of more flexible, data-driven approaches to high-dimensional DA that can leverage non-Gaussian priors without suffering from the same curse of dimensionality as particle filters.
@@
Recent progress in offline inverse problems—such as those in medical imaging, astrophysics, and computer vision—has shown that flow- and diffusion-based generative models can serve as powerful priors for conditional sampling (see, e.g., Zheng et al. (2025); Chung et al. (2025); Zhao et al. (2025) for a recent benchmark). However, extending these ideas to sequential filtering presents unique challenges. One possibility is to learn a generative prior for the predictive distribution $p({\bm{x}}_{n}|{\bm{y}}_{1:n-1})$, but this would require retraining the model at every time step $n$ (Bao et al., 2024a), making it impractical for operational use. Another strategy is to use a pre-trained generative forecast model to guide towards observations (Chen et al., 2025; Savary et al., 2025), though such models require the use of specialized dynamical models. Alternatively, approaches based on approximating the smoothing distribution $p({\bm{x}}_{0:n}|{\bm{y}}_{1:n})$ (Rozet and Louppe, 2023) exist. However, they are memory-intensive and scale poorly with the number of time steps. We elaborate on how these approaches relate to our proposed method in Section 3.3.
@@ 1.1 Contributions
In this work, we present DAISI (Data Assimilation with Inverse sampling using Stochastic Interpolants), a scalable filtering algorithm built on flow-based generative models. DAISI avoids retraining at each assimilation step by leveraging a stationary pre-trained generative prior to condition on new observations via guidance. To integrate dynamical information, DAISI couples this generative prior with a forecast model that advances an ensemble of states in time. Following the forecast, an inverse-sampling step runs the generative SDE backward from the forecast ensemble, mapping forecasted states to latent variables. These latent states then serve as initial conditions for conditional sampling under the learned prior. The full procedure is illustrated in Figure 1.
@@
    Modular design, supporting any flow-based generative model and gradient-based guidance method.
@@
    Expressive uncertainty quantification, capturing complex, multimodal, high-dimensional posteriors under sparse, noisy, and nonlinear observations.
@@ 2.1 Data Assimilation
where $\mathcal{F}_{\sharp}\pi$ denotes the pushforward of the measure $\pi$ with respect to a map $\mathcal{F}$, defined by $\mathcal{F}_{\sharp}\pi:=\mathrm{Law}(\mathcal{F}(X))$ for $X\sim\pi$. Starting from $\pi_{0}$, repeated applications of (3)–(4) yields the sequence $\pi_{1},\ldots,\pi_{N}$. Although this recursion is well defined, computing these measures exactly is typically infeasible in practice. Our aim is therefore to develop a practical and scalable approximation to the cycle (3)–(4), capable of operating reliably in high-dimensional settings with sparse, nonlinear observations.
@@ 2.2 Stochastic Interpolants for Generative Modeling
Our proposed DA method leverages recent advances in generative modeling. In the static setting (i.e., ignoring time evolution), flow-based models enable sampling from a target distribution $\rho_{1}$ by learning a transport map from a simple latent distribution $\rho_{0}$. Such transports can be parameterized in multiple ways; in this work, we adopt the framework of stochastic interpolants (Albergo et al., 2023), though closely related formulations appear in the probability flow ODE for diffusion models (Song et al., 2021) and in flow matching (Lipman et al., 2023).
@@
Let $\rho_{t}:=\mathrm{Law}({\bm{z}}_{t})$ denote the distribution of ${\bm{z}}_{t}$ with density $p_{t}$, and define its score ${\bm{s}}(t,{\bm{z}}):=\nabla\log p_{t}({\bm{z}})$ and drift ${\bm{b}}(t,{\bm{z}}):=\mathbb{E}[\dot{{\bm{z}}}_{t}|{\bm{z}}_{t}={\bm{z}}]$. Albergo et al. (2023) show that for any non-negative $\epsilon\in C^{0}([0,1])$, the SDE
@@ 2.3 Conditional Generation via Guidance
In general, the guidance term is intractable as we usually only have access to $p({\bm{y}}|{\bm{z}}_{1})$ but not $p({\bm{y}}|{\bm{z}}_{t})=\mathbb{E}_{{\bm{z}}_{1}|{\bm{z}}_{t}}[p({\bm{y}}|{\bm{z}}_{1})]$. Various approaches exist to approximate this term, e.g. diffusion posterior sampling (DPS) (Chung et al., 2023) and moment matching posterior sampling (MMPS) (Rozet et al., 2024) (see Appendix A.5 for details and (Daras et al., 2024) for a survey on guidance methods).
@@ 3 Method
Our method, DAISI, performs filtering by combining the strengths of ensemble-based methods with the expressive priors offered by flow-based generative models. This is conceptually similar to the hybrid ensemble-variational (EnVar) method in classical DA (Hamill and Snyder, 2000; Lorenc, 2003), which blend a static climatological background covariance $\mathbf{P}$ with an “error-of-the-day” ensemble covariance to represent forecast uncertainty. This compensates for the limitations of either component alone: relying solely on a static background covariance ignores the dynamical evolution of errors, while using only a small ensemble fails to characterize high-dimensional uncertainties accurately.
DAISI builds on this philosophy but moves beyond Gaussian background covariances $\mathbf{P}$ used in hybrid EnVar by considering a full background measure $\mathbb{P}_{\infty}$, learned using flow-based generative models. In practice, we take $\mathbb{P}_{\infty}$ to be the invariant measure of the dynamical system (1), whose samples can be approximated from trajectories. While its existence and ergodicity are not generally guaranteed, it is a reasonable assumption for many systems, including geophysical flows. Given the learned prior $\mathbb{P}_{\infty}$, we can in principle assimilate the observation ${\bm{y}}_{n}$ by sampling from the posterior
@@
To achieve this, we introduce inverse sampling, which transfers dynamical information from the ensemble forecast into the latent space of the generative model. Given forecast particles $\{\hat{{\bm{x}}}^{(j)}_{n}\}_{j=1}^{J}$, this proceeds by solving the unconditional backward SDE (7) from $t=1$ to $t=t_{\min}\in(0,1)$, using each $\hat{{\bm{x}}}_{n}^{(j)}$ as terminal conditions. The resulting latent variables $\{{\bm{z}}_{t_{\min},n}^{(j)}\}_{j=1}^{J}$ encode the forecast information in “noise space”. We can then perform conditional sampling by integrating the guided forward SDE (8) from $t_{\min}$ to 1, initialized at these latent states. This produces updated particles $\{{\bm{x}}_{n}^{(j)}\}_{j=1}^{J}$ that are samples of $\mathbb{P}_{\infty}^{\bm{y}}$, while simultaneously containing information about the ensemble $\{\hat{{\bm{x}}}^{(j)}_{n}\}_{j=1}^{J}$ through its latent representations $\{{\bm{z}}_{t_{\min},n}^{(j)}\}_{j=1}^{J}$.
@@ 3.1 Does DAISI sample from the filtering distribution?
where $\hat{\pi}_{n}$ denotes the predictive distribution at time $n$, and the expectation is over the Brownian motions used when integrating the backward and forward SDEs. For DAISI to function as a reliable filter, this measure should approximate the true filtering distribution $\pi_{n}$ closely. While there is no reason to believe that (13) matches $\pi_{n}$ exactly, we argue that we can achieve a good approximation by tuning the hyperparameters $t_{\min}$ and $\epsilon$ appropriately.
@@
For simplicity, assume that $\epsilon=0$ and introduce the shorthands $\pi_{n}^{\text{DAISI}}:=\pi_{n,0,0}^{\text{DAISI}}$, $\Phi:=\Phi_{0,1}^{0}$ and $\Psi:=\Psi_{1,0}^{0}$. Similarly, write $\Phi^{\bm{y}}:=\Phi_{0,1}^{{\bm{y}},0}$ and $\Psi^{\bm{y}}:=\Psi_{1,0}^{{\bm{y}},0}$. Note that in this case, we have $\Psi=\Phi^{-1}$ and $\Psi^{\bm{y}}=(\Phi^{\bm{y}})^{-1}$. To understand the effect of $t_{\min}$, we consider the limiting cases $t_{\min}\rightarrow 1$ and $t_{\min}\rightarrow 0$. By time-continuity of the flows on $[0,1]$, the first limit $t_{\min}\rightarrow 1$ trivially yields $\pi_{n,t_{\min},0}^{\text{DAISI}}\rightarrow\hat{\pi}_{n}\propto f_{n}\mathbb{P}_{\infty}$. Comparing with (16), we observe that although the density ratios agree, the underlying base measures differ.
@@
$\widetilde{\rho}_{0}(\mathrm{d}{\bm{x}}):=\Psi_{\sharp}\hat{\pi}_{n}(\mathrm{d}{\bm{x}})$ (17)
$\propto f_{n}(\Phi({\bm{x}}))\Psi_{\sharp}\mathbb{P}_{\infty}(\mathrm{d}{\bm{x}})=f_{n}(\Phi({\bm{x}}))\rho_{0}(\mathrm{d}{\bm{x}}),$ (18)
where we used that $\Psi_{\sharp}\mathbb{P}_{\infty}=\rho_{0}$. Next applying the guided flow in the second step gives us
$\pi_{n}^{\text{DAISI}}(\mathrm{d}{\bm{x}})=\Phi^{{\bm{y}}}_{\sharp}\widetilde{\rho}_{0}(\mathrm{d}{\bm{x}})$ (19)
@@
where $g_{n}^{{\bm{y}}}({\bm{x}}):=f_{n}(\Phi(\Psi^{\bm{y}}({\bm{x}})))$ and we used that $\Phi^{{\bm{y}}}$ transports $\rho_{0}$ to $\mathbb{P}_{\infty}^{\bm{y}}$ by construction. Comparing (20) with (16), we see that in this limit, the base measures match, while the density ratios differ. Thus, the extremes $t_{\min}\rightarrow 1$ and $t_{\min}\rightarrow 0$ each match only one component of the true filtering distribution, suggesting the existence of an intermediate $t_{\min}\in(0,1)$ that provides the best trade-off between these two types of mismatch.
In Figure 2, we illustrate this trade-off using a simple 1D toy problem (see Appendix D.2 for details). When $t_{\min}=0.01$, the distribution $\pi^{\text{DAISI}}_{n}$ (in orange) shares characteristics of $\mathbb{P}^{\bm{y}}_{\infty}$, as predicted from (20). Conversely, for a large value such as $t_{\min}=0.6$, the distribution closely matches $\hat{\pi}_{n}$, as expected. The intermediate setting $t_{\min}=0.3$ yields a distribution that interpolates between these two extremes, yielding a distribution that better aligns with the filtering distribution, which combines features of both $\mathbb{P}^{\bm{y}}_{\infty}$ and $\hat{\pi}_{n}$ — as reflected by the noticeably lower Maximum Mean Discrepancy (MMD) score (Gretton et al., 2012).
@@
Assume that the marginal laws $\{\rho_{t}\}_{t\in[0,1]}$ and $\{\rho_{t}^{\bm{y}}\}_{t\in[0,1]}$ of the interpolants bridging $\rho_{0}=\mathcal{N}(\bm{0},\mathbf{I})$ with $\rho_{1}=\mathbb{P}_{\infty}$ and $\rho_{1}^{\bm{y}}=\mathbb{P}_{\infty}^{\bm{y}}$, respectively, satisfy a log-Sobolev inequality with constant $\lambda>0$, uniform in $t$. Then,
@@
In particular, when $\epsilon\rightarrow\infty$, we have $\pi_{n,0,\epsilon}^{\text{DAISI}}\rightarrow\mathbb{P}_{\infty}^{\bm{y}}$ and therefore all ensemble information is lost. In practice, we find that taking a small but non-zero $\epsilon$ is useful for mitigating the tendency of particles from collapsing to a single state as assimilation progresses. In our 1D toy experiment (Appendix D.2), we show that tuning both $t_{\min}$ and $\epsilon$ yields better performance than adjusting $t_{\min}$ alone.